Royal Lawncare Company produces and sells two packaged products—Weedban and Greengrow. Revenue and cost information relating to the products follow:

Product

Weedban Greengrow
Selling price per unit $ 12.00 $ 33.00
Variable expenses per unit $ 2.70 $ 12.00
Traceable fixed expenses per year $ 137,000 $ 43,000

Common fixed expenses in the company total $97,000 annually. Last year the company produced and sold 41,000 units of Weedban and 20,500 units of Greengrow.

Required:

Prepare a contribution format income statement segmented by product lines.
Answer #1

E 1 2 Sales 3 Variable expenses 4 Contribution margin 5 Traceable fixed expense 6 Segment margin 7 Common Fixed expense 8 NetA B. 1 Total company Weedban = 41000 Units Greengrow = 20500 Units 2 Sales =C2+D2 =41000*12 =20500*33 3 Variable expenses =D3